OpenStudy (anonymous):

The sum of three consecutive odd integers is 207. What are the integers? The first integer is The second integer is The third integer is

OpenStudy (yttrium):

x+1+(x+3)+(x+5)=207 3x = 198 x = 66 x+1 = 67 x+3 = 69 x+5 = 71